OpenStudy (anonymous):

Personal Consumption Although there was an economic downturn in 2008, personal consumption has grown overall. Table 11.15 lists (approximate) personal consumption C in dollars for selected years x. TABLE 11.15 Personal Consumption (dollars) x 1959 1982 1990 1998 2009 C $300 $2000 $4000 $6000 $10,000 Source: BEA. (a)Make a scatterplot of these data. Use the window [1955, 2010, 10] by [0, 12000, 2000].* (b) Find a quadratic function given by C(x) = a(x - 1959)^2 + k (x) = 3.9(x - 1959)2 + 300 that models the data. Graph C and the data. (Answers may vary.)* (c) Estimate when personal consumption was $8000. Compare your answer to the actual value of 2005. About 2004; 1 year too soon (d) Use your function C to predict consumption in 2020 to the nearest thousand dollars.

OpenStudy (anonymous):

[C(x)=a(x-1959)^2+K] a=3.21 K=300 (1959,300)(1982,2000)(1990,4000)(1998,6000)(2009,10000)

OpenStudy (anonymous):

Can anyone help with this?

OpenStudy (anonymous):

x 1959, 1982, 1990, 1998, 2009 C $300, $2000, $4000, $6000, $10000 C(1959)=a(1959-1959)^2+K 300=a(0)^2+K 300=K C(1982)=a(1982-1959)^2+300 2000=a(31)^2+300 2000=a(529)+300 1700=529a 3.21=a

OpenStudy (anonymous):

Personal Consumption Although there was an economic downturn in 2008, personal consumption has grown overall. Table 11.15 lists (approximate) personal consumption C in dollars for selected years x. TABLE 11.15 Personal Consumption (dollars) x 1959 1982 1990 1998 2009 C $300 $2000 $4000 $6000 $10,000 Source: BEA. (a)Make a scatterplot of these data. Use the window [1955, 2010, 10] by [0, 12000, 2000].* (b) Find a quadratic function given by C(x) = a(x - 1959)^2 + k (x) = 3.9(x - 1959)2 + 300 that models the data. Graph C and the data. (Answers may vary.)* (c) Estimate when personal consumption was $8000. Compare your answer to the actual value of 2005. About 2004; 1 year too soon (d) Use your function C to predict consumption in 2020 to the nearest thousand dollars.

OpenStudy (anonymous):

@zepdrix can you help?

OpenStudy (anonymous):

This is wearing me out.

OpenStudy (anonymous):

me too! lol

OpenStudy (anonymous):

The crazy thing is that the word document I am coping from doesn't show b, with the (x)=3.9(x-1959)^2+K.

OpenStudy (anonymous):

I have know idea where that is coming from.

OpenStudy (anonymous):

I solved for K like this C(x)=a(x-1959)^2+K 300=a(1959-1959)^2+K 300 =a(0)^2+K 300=K

OpenStudy (anonymous):

I used K to find a C(x)=a(x-1959)^2+K C(1982)=a(1982-1959)^2+K 2000=a(23)^2+300 2000=a(529)+300 2000=529a+300 2000-300=529a+300-300 1700=529a 3.21=a